# FuSeFL: Fully Secure and Scalable Cross-Silo Federated Learning

### v2

@@ Abstract
Federated Learning (FL) enables collaborative model training without centralizing client data, making it attractive for privacy-sensitive domains. While existing approaches employ cryptographic techniques such as homomorphic encryption, differential privacy, or secure multiparty computation to mitigate inference attacks—including model inversion, membership inference, and gradient leakage—they often suffer from high computational, communication, or memory overheads. Moreover, many methods overlook the confidentiality of the global model itself, which may be proprietary and sensitive. These challenges limit the practicality of secure FL, especially in cross-silo deployments involving large datasets and strict compliance requirements.
We present FuSeFL, a fully secure and scalable FL scheme designed for cross-silo settings. FuSeFL decentralizes training across client pairs using lightweight secure multiparty computation (MPC), while confining the server’s role to secure aggregation. This design eliminates server bottlenecks, avoids data offloading, and preserves full confidentiality of data, model, and updates throughout training. FuSeFL defends against inference threats, achieves up to 95% lower communication latency and 50% lower server memory usage, and improves accuracy over prior secure FL solutions, demonstrating strong security and efficiency at scale.
@@ I Introduction
Federated Learning (FL) is a distributed machine learning paradigm that enables multiple clients to collaboratively train a global model without sharing their local data [1, 2, 3, 4, 5, 6, 7]. This paradigm is particularly suited for scenarios where legal, privacy, or infrastructural constraints prevent centralized data collection [8, 5]. In FL, a central server distributes a model to the clients, who perform local training and return updates, such as gradients or model weights. These updates are aggregated to refine the global model, all while the raw data remains on client devices, preserving confidentiality and reducing transmission costs. While FL offers clear advantages in preserving data locality, it remains vulnerable to indirect privacy breaches. In particular, submitted gradients or model updates can unintentionally encode patterns from the underlying data, enabling adversaries to extract sensitive information from these values [9, 4, 10, 11].
@@
To mitigate such risks, a wide range of techniques have been proposed to protect the privacy of clients’ data during training. Homomorphic encryption (HE) enables secure aggregation over encrypted updates without decryption [3, 4], but incurs high computational and memory costs, often slowing down training by several orders of magnitude [12]. Differential privacy (DP) defends against inference attacks by adding noise to updates [5, 6], though stronger privacy budgets degrade model accuracy [12]. Secure multi-party computation (MPC) allows collaborative aggregation without exposing individual inputs [7, 13, 11, 14, 4, 15], but suffers from growing communication overhead as the number of clients increases [12].
@@
Beyond data privacy, a critical yet often underappreciated concern in federated learning is the privacy of the model itself, particularly from the perspective of the model owner. In typical FL deployments, the global model, often developed through significant intellectual and financial investment, is shared with all participating clients, some of whom may be semi-honest or untrusted [16]. This exposes the model to risks such as extraction, reverse engineering, or unauthorized use. The concern becomes even more pressing in cross-silo FL, where clients are typically large institutions, such as hospitals, banks, or research labs, that operate under strict privacy regulations and contractual obligations. In such settings, the model is not only a valuable proprietary asset but also a potential vector for secondary leakage of sensitive information. For example, attackers may perform model inversion to reconstruct private inputs or launch membership inference attacks to determine whether a particular data point was part of the training set [17]. Ensuring model confidentiality is therefore essential, not only to protect intellectual property but also to comply with legal, regulatory, and contractual boundaries.
Several recent works have attempted to address model privacy in federated learning, though each approach faces significant limitations that hinder its practicality, particularly in cross-silo deployments where efficiency, scalability, and regulatory compliance are critical.
Watermarking-based techniques have been proposed to address model ownership and intellectual property protection [18, 19]. These methods embed verifiable patterns, such as trigger inputs or parameter, level watermarks—into the model, allowing the owner to later demonstrate unauthorized use. However, watermarking does not prevent theft or misuse during training; it only enables detection after the fact. In domains like healthcare and finance, where compliance and security must be enforced upfront, such reactive mechanisms are insufficient.
@@
WW-FL [20] adopts a fully offloaded training architecture in which secure multi-party computation (MPC) is performed across distributed clusters of servers. While this design offers strong security guarantees, it requires clients to upload their entire datasets to third-party servers, contradicting the foundational principle of cross-silo FL, where data should remain local to each silo. In addition, this architecture introduces a trade-off between communication overhead, memory usage, and the number of server clusters. When dealing with thousands of clients, increasing the number of servers raises the system’s cost while reducing communication time and memory usage. However, reducing the number of server clusters leads to significant communication overhead, as large volumes of sensitive data must be transmitted over the network, increasing both latency and memory consumption. Moreover, WW-FL depends on geographically distributed compute infrastructure, which elevates deployment complexity and operational costs. In cross-silo environments, where clients such as hospitals or financial institutions already possess rich compute and storage capabilities, outsourcing training to external servers is not only inefficient but also unnecessary.
@@
A similar line of work proposes offloading both the model and client data to centralized secure servers that perform training using cryptographic techniques [21] such as MPC [22, 23, 24, 25, 26, 27]. While these systems reduce the trust placed in individual clients, they impose substantial communication burdens due to repeated transfers of large training datasets. This issue is exacerbated in cross-silo FL, where each client holds sizable, high-value data. Additionally, centralizing sensitive datasets on untrusted servers risks violating regulatory constraints, institutional data policies, and contractual agreements, making these approaches ill-suited for regulated domains.
@@
AriaNN [25] introduces a hybrid federated scheme that jointly protects model and data privacy by secret-sharing the model between the clients and server and distributing the computation using MPC. This design prevents either party from observing the full model or training data in plaintext. While more aligned with FL principles, AriaNN suffers from serious scalability limitations. Each client must train its share of the model in coordination with the server, typically in a sequential fashion due to server-side compute constraints. Communication overhead grows linearly with client count; even modest datasets like MNIST can require tens of megabytes per client, and this cost becomes prohibitive for large datasets common in siloed environments. Furthermore, the server must maintain a dedicated model instance per client in addition to the global model, leading to unsustainable memory overhead for large-scale models such as VGG-16.
In summary, existing federated learning approaches for cross-silo settings struggle to scale beyond a few tens of clients due to excessive computational, communication, and memory overheads. Techniques like homomorphic encryption and secure model-sharing, while effective in preserving privacy, impose severe performance costs. In particular, efforts to simultaneously protect both data and model confidentiality exacerbate these challenges, rendering current solutions impractical for real-world, large-scale deployments.
To address the limitations of prior work, we propose FuSeFL, a fully secure and scalable federated learning framework tailored for cross-silo environments. FuSeFL overcomes the limitations of prior work by ensuring end-to-end confidentiality of both the data and the model, while also scaling efficiently to large client populations. The key idea is to decentralize training across client groups while confining the server’s role to lightweight secure aggregation.
@@
In FuSeFL, clients are dynamically paired to form collaborative groups. Within each group, the global model is secret-shared between the two clients, and training is jointly executed using secure multi-party computation (MPC) over their secret-shared local data. This structure prevents any party from accessing the model or data in plaintext. After training, each group submits secret-shared updates to two non-colluding aggregation servers, which perform secure aggregation without learning individual contributions.
This design resolves the critical issues in prior solutions: it avoids server-side bottlenecks, eliminates the need for data offloading, and reduces communication overhead. Moreover, by keeping both the model and updates secret-shared throughout training and aggregation, FuSeFL protects against gradient leakage, model update inference, membership inference, and model inversion attacks.
Our evaluation shows that FuSeFL (1) reduces server memory usage by nearly 50%, (2) maintains nearly constant training time as the number of clients increases, achieving linear throughput scalability, and (3) cuts communication latency by 95% compared to the state-of-the-art schemes in large-scale deployments. FuSeFL also improves model accuracy by up to 1.1% over the federated adaptation of AriaNN on MNIST.
@@
  • A fully secure FL scheme: We develop FuSeFL, a federated learning scheme that holistically secures model parameters, client data, and intermediate updates using lightweight MPC across client groups, addressing both model and data privacy within a unified design (§V-A).
  • Decentralized and scalable training: FuSeFL eliminates server-side training bottlenecks by distributing computation among client pairs. This design supports parallelism, avoids reliance on third-party compute infrastructure, and is well-suited for cross-silo FL environments with powerful clients (§VII-B).
  • Practical efficiency with robust privacy guarantees: Our approach significantly reduces memory and communication overheads while mitigating key inference threats such as model inversion, gradient leakage, and membership inference, making FuSeFL scalable to real-world deployments involving many silos (§VII-D, §VII-C).
@@ II-A Limitations of Existing Work
Despite increased attention to privacy in federated learning, fully secure FL remains constrained by fundamental system bottlenecks. We identify three key limitations:
@@
    Model Exposure and Trust Assumptions: Many existing FL schemes expose the full global model to clients during training or inference, risking model theft or misuse, particularly in regulated domains where model intellectual property is sensitive [28, 18, 19]. Moreover, several privacy-preserving methods, such as WW-FL[20], offload client data or models to trusted third-party servers using MPC or HE protocols [22, 23, 24, 25, 29, 30]. These trust assumptions violate the core principle of cross-silo FL, which requires data to remain within institutional boundaries [8, 31].
@@
    Communication and Storage Overheads: Secure training introduces high communication and memory demands. For example, training Network1 over MNIST with AriaNN-FL requires more than 10 MB of data exchange per client per epoch [25]. With hundreds of clients, this quickly overwhelms bandwidth limits. Additionally, the server must retain a copy of each client’s model update until aggregation is complete. For large models like VGG-16, this results in excessive memory usage: as illustrated in Fig. 1 (right axis), AriaNN-FL stores 128 full model instances when training with 128 clients. In schemes such as WW-FL, where the model and user data are offloaded to third-party servers, communication latency becomes significant when submitting user data to the training server cluster, as the number of server clusters is limited due to their associated cost.
Figure 1: Comparison of AriaNN-FL with FuSeFL, (Left vertical axis) in terms of secure training time for one epoch and (Right vertical axis) in terms of the number of model instances stored in one aggregator. All these experiments use 16 parallel computing cores for a varying number of clients. By increasing the number of clients, the training time and the number of model instances increase linearly in AriaNN-FL.
@@ II-B Our Goals
FuSeFL aims to deliver a practical, fully secure, and scalable cross-silo federated learning framework that safeguards both user data and the model itself, an often-overlooked dimension in secure FL literature. Designed for real-world cross-silo deployments with potentially several hundred or even thousands of clients, FuSeFL balances strict privacy guarantees with system-level scalability and efficiency. Specifically, it seeks to:
  • Ensure end-to-end confidentiality of data and models: Unlike conventional FL, where clients have full access to the global model, FuSeFL guarantees that both the model and its updates remain secret-shared and are never revealed in plaintext to any single party.
  • Remove server-side computational bottlenecks: By decentralizing local training to client groups, FuSeFL avoids the per-client training load imposed on servers in frameworks like AriaNN-FL. Servers are instead limited to lightweight secure aggregation, improving scalability and efficiency.
  • Reduce communication and memory overhead: FuSeFL forms client pairs that collaboratively train and send a single secret-shared model update per group, rather than having each client transmit updates independently. This design cuts communication volume and nearly halves server-side memory usage, improving resource efficiency in large-scale deployments.
@@ III Threat Model and Security Considerations
Deployment model and trust assumptions: FuSeFL targets scalable cross-silo federated learning where each client, such as a hospital or financial institution, holds a distinct set of samples over a shared feature space, a concept known as horizontal federated learning [32]. Clients are governed by regulatory frameworks and formal agreements such as HIPAA or GDPR, which limit the risk of deviation from the protocol.
The trust model assumes: (1) Clients do not collude with others in the same training group. Each may attempt to infer private information from its own view but adheres to the protocol. (2) Aggregation servers are honest-but-curious, do not collude with one another or with clients, and jointly perform MPC-based aggregation. (3) The model owner is trusted to manage protocol coordination but cannot access client data or intermediate model states.
To reduce the risk of collusion and information leakage across rounds, FuSeFL employs dynamic client grouping, reassigning clients into new training groups in each round to prevent repeated interactions. While our protocol assumes no collusion within a group during a single round, this is a well-founded assumption in cross-silo federated learning[8, 31], where clients are typically autonomous organizations with independent data silos, isolated infrastructure, and legally enforced compliance boundaries. Such structural and regulatory separation significantly reduces the feasibility of coordinated behavior required for collusion. FuSeFL is architected to support threshold secret sharing as a drop-in extension, enabling privacy preservation even with up to $t$ colluding clients per group. We plan to implement and evaluate this enhancement in future work.
Threat model: We consider a semi-honest (honest-but-curious) adversary who follows the protocol but may attempt to infer private information from observed data. The system involves three types of participants: clients, aggregation servers, and a model owner. Clients perform local training using secret-shared model parameters. Aggregation servers coordinate secure aggregation without learning the underlying values. The model owner is responsible for initializing the global model and orchestrating training (as shown in Fig. 2).
@@
These guarantees mitigate risks such as unauthorized observation, gradient leakage, and reconstruction attacks. Since the model is never exposed in plaintext outside the trusted boundary, no untrusted party can access it to carry out membership inference or model inversion attacks aimed at extracting information about the training user data. FuSeFL enforces these protections using efficient secure MPC protocols.
Out-of-scope threats: FuSeFL does not currently defend against poisoning attacks [33] and malicious participants who deviate from the protocol, submit corrupted updates, or act in a Byzantine manner. Future extensions may incorporate robust aggregation, verifiable computation, or trusted execution to address these threats. Side-channel attacks[34, 35, 36] such as timing[34] or traffic analysis[36] are also out of scope and would require separate system-level countermeasures.
Figure 2: Overview of the threat model. Each party retains private data within its trusted boundary. Adversaries may attempt to infer information about the data or model from the untrusted boundary.

### v3

# FuSeFL: Fully Secure and Scalable Federated Learning

@@ Abstract
Federated Learning (FL) enables collaborative model training without centralizing client data, making it attractive for privacy-sensitive domains. While existing approaches employ cryptographic techniques such as homomorphic encryption, differential privacy, or secure multiparty computation to mitigate inference attacks, including model inversion, membership inference, and gradient leakage, they often suffer from high computational and memory overheads. Moreover, many methods overlook the confidentiality of the global model itself, which may be proprietary and sensitive. These challenges limit the practicality of secure FL, especially in settings that involve large datasets and strict compliance requirements.
We present FuSeFL, a Fully Secure and scalable FL scheme, which decentralizes training across client pairs using lightweight MPC, while confining the server’s role to secure aggregation, client pairing, and routing. This design eliminates server bottlenecks, avoids full data offloading, and preserves full confidentiality of data, model, and updates throughout training. Based on our experiment, FuSeFL defends against unauthorized observation, reconstruction attacks, and inference attacks such as gradient leakage, membership inference, and inversion attacks, while achieving up to $13\times$ speedup in training time and 50% lower server memory usage compared to our baseline.
@@ I Introduction
Federated Learning (FL) is a distributed machine learning paradigm that enables multiple clients to collaboratively train a global model without sharing their local data [1, 2, 3, 4, 5, 6, 7]. This paradigm is particularly suited for scenarios where legal, privacy, or infrastructural constraints prevent centralized data collection [8, 5]. In FL, a central server distributes a model to the clients, who perform local training and return updates, such as gradients or model weights. These updates are aggregated to refine the global model, all while the raw data remains on client devices, preserving confidentiality and reducing transmission costs. While FL offers clear advantages in preserving data locality, it remains vulnerable to significant privacy breaches, where (1) submitted gradients or model updates can unintentionally encode patterns from the underlying data, allowing adversaries to extract sensitive information [9, 4, 10, 11], and (2) attackers may attempt to steal or misuse the offloaded model for malicious or competitive gain.
@@
To mitigate users’ data privacy concerns, a wide range of techniques have been proposed to protect the privacy of clients’ data during training. Homomorphic encryption (HE) enables secure aggregation over encrypted updates without decryption [3, 4], but incurs high computational and memory costs, often slowing down training by several orders of magnitude [12]. Differential privacy (DP) defends against inference attacks by adding noise to updates [5, 6], though stronger privacy budgets degrade model accuracy [12]. Secure multi-party computation (MPC) allows collaborative aggregation without exposing individual inputs [7, 13, 11, 14, 4, 15], but communication overhead grows with the number of clients [12].
@@
Beyond data privacy, a critical yet often underappreciated concern in FL is the privacy of the model itself, particularly from the perspective of the model owner. In typical FL deployments, the global model, often developed through significant intellectual and financial investment, is shared with all participating clients, some of whom may be semi-honest or untrusted [16]. This exposes the model to risks such as extraction, reverse engineering, or unauthorized use. The concern becomes even more pressing when clients are typically large institutions, such as hospitals, banks, or research labs, that operate under strict privacy regulations and contractual obligations. In such settings, the model is not only a valuable proprietary asset but also a potential vector for secondary leakage of sensitive information. For example, attackers may perform model inversion to reconstruct private inputs or launch membership inference attacks to determine whether a particular data was part of the training set [17]. Ensuring model confidentiality is therefore essential, not only to protect intellectual property but also to comply with legal, regulatory, and contractual boundaries. Several recent works have attempted to address model privacy in FL[18, 19, 20, 21, 22, 23, 24], though each approach faces significant limitations that hinder its practicality, particularly in settings where efficiency, scalability, and regulatory compliance are critical. Watermarking-based techniques have been proposed to address model ownership and intellectual property protection [25, 26]. These methods embed verifiable patterns, such as trigger inputs or parameter-level watermarks, into the model, allowing the owner to later demonstrate unauthorized use. However, watermarking does not prevent theft or misuse during training; it only enables detection after the fact. In domains like healthcare and finance, where compliance and security must be enforced upfront, such reactive mechanisms are insufficient.
@@
WW-FL [18] adopts a fully offloaded training architecture in which MPC is performed across distributed clusters of servers. While this design offers strong security guarantees, it requires clients to upload their entire datasets to third-party servers, contradicting the foundational principle of FL, where data should remain local to each client. In addition, this architecture introduces a trade-off between communication overhead, memory usage, and the number of server clusters. When dealing with thousands of clients, increasing the number of servers raises the system’s cost while reducing communication time and memory usage.
Moreover, WW-FL depends on geographically distributed compute infrastructure, which elevates deployment complexity and operational costs. In addition, in settings such as cross-silo environments, where clients (hospitals or financial institutions) already possess rich compute and storage capabilities, outsourcing training to external servers is not only inefficient but also unnecessary.
@@
A similar line of work proposes offloading both the model and client data to centralized secure servers that perform training using cryptographic techniques [27] such as MPC [19, 20, 21, 22, 23, 24]. While these systems reduce the trust placed in individual clients, they impose substantial communication burdens due to repeated transfers of large training datasets. Additionally, centralizing sensitive datasets on untrusted servers risks violating regulatory constraints, institutional data policies, and contractual agreements, making these approaches ill-suited for regulated domains.
@@
AriaNN [22] introduces a hybrid FL scheme that jointly protects model and data privacy by secret-sharing the model between the clients and server using MPC. This design prevents either party from observing the full model or training data in plaintext. While more aligned with FL principles, AriaNN suffers from serious scalability limitations. Each client must train its share of the model in coordination with the server, typically in a sequential fashion due to server-side compute constraints. Furthermore, the server must maintain a dedicated model instance per client in addition to the global model, leading to unsustainable memory overhead for large-scale models such as VGG-16. In summary, existing fully secure FL approaches struggle to scale beyond a few tens of clients due to excessive server-side computational and memory overheads. Techniques like HE and secure model sharing, while effective in preserving privacy, impose severe performance costs. In particular, efforts to simultaneously protect both data and model confidentiality exacerbate these challenges, rendering current solutions impractical for real-world, large-scale deployments.
@@
To address the limitations of prior work, we propose FuSeFL, a fully secure and scalable federated learning (FL) framework. FuSeFL ensures end-to-end confidentiality of both data and models while efficiently scaling to tens of thousands of clients without overloading the server. Its core idea is to decentralize training across dynamically formed client groups, while confining the server’s role to lightweight secure aggregation, client pairing, and data routing. In FuSeFL, clients are dynamically paired to form collaborative groups, where the global model is secret-shared between the two clients, and training is jointly executed using MPC over their secret-shared local data. This setup prevents any participant from accessing data or model parameters in plaintext. After local training, each client pair submits secret-shared updates to two servers, which perform secure aggregation without learning any private information.
[…]
@@
  • A fully secure FL scheme: We develop FuSeFL, a FL scheme that holistically secures model parameters, client data, and intermediate updates using lightweight MPC across client groups, addressing both model and data privacy within a unified design (§V-A).
  • Decentralized and scalable training: FuSeFL eliminates server-side training bottlenecks by distributing heavy computations among client pairs, and it only relies on the server for performing lightweight computations (§VII-A).
  • Practical efficiency with robust privacy guarantees: Our approach significantly reduces memory and computation overheads while mitigating key inference threats such as model inversion, gradient leakage, and membership inference, making FuSeFL scalable to real-world deployments involving many clients (§VII-C, §VII-A).
@@ II-A Limitations of Existing Work
Despite increased attention to privacy in FL, fully secure FL remains constrained by fundamental system bottlenecks. We identify three key limitations:
@@
    Model Exposure and Trust Assumptions: Many FL schemes expose the global model to clients during training, risking model theft or misuse, particularly in regulated domains where model is sensitive [28, 25, 26]. Moreover, several privacy-preserving methods, such as WW-FL[18], offload client data or models to trusted third-party servers using MPC or HE protocols [19, 20, 21, 22, 29, 30] while assuming that the computing parties do not collude. These trust assumptions are not realistic or violate the core principle of FL, which requires data to remain within clients’ boundaries [8, 31].
@@
    Storage Overheads: Secure training introduces high memory demands. This is because the server must retain a copy of each client’s model update until aggregation is complete. For large models like VGG-16, this results in excessive memory usage: as illustrated in Fig. 1 (right axis), AriaNN-FL stores 128 full model instances when training with 128 clients.
Figure 1: AriaNN-FL vs. FuSeFL, (Left vertical axis) in terms of secure training time for one epoch and (Right vertical axis) in terms of the number of model instances stored in one aggregator.
@@ II-B Our Goals
FuSeFL aims to deliver a practical, fully secure, and scalable FL framework that safeguards both user data and the model itself, an often-overlooked dimension in secure FL literature. Designed for real-world deployments with potentially several thousands of clients, FuSeFL balances strict privacy guarantees with system-level scalability and efficiency. Specifically, FuSeFL seeks to:
  • Ensure end-to-end confidentiality of data and models: Unlike conventional FL, where clients have full access to the global model, FuSeFL guarantees that the model and its updates remain secret-shared and are never revealed in plaintext to untrusted parties.
  • Remove server-side computational bottlenecks: By decentralizing local training to client groups, FuSeFL avoids per-client training load imposed on servers in frameworks like AriaNN-FL. Servers’ computations are instead limited to lightweight secure aggregation and client pairing, improving scalability and efficiency.
  • Reduce memory overhead: FuSeFL forms client pairs that collaboratively train and send a single secret-shared model update per group, rather than having each client transmit updates independently. This design nearly halves server-side memory usage, improving resource efficiency in large-scale deployments.
@@ III Threat Model and Security Considerations
Deployment Model and Trust Assumptions: FuSeFL targets scalable FL where each client, such as a hospital, financial institution, or mobile device, holds a distinct set of samples over a shared feature space, i.e., horizontal FL [32]. Clients are governed by regulatory frameworks and formal agreements such as HIPAA or GDPR, which limit the risk of deviation from the protocol.
The trust model assumes: (1) Clients do not know whom they are pairing with, as their communication is routed through a trusted party. Due to this anonymity, clients in the same group cannot collude with each other or with the servers. Collusion among clients outside each training pair is not problematic, since the shares they hold do not allow the reconstruction of sensitive information. Clients are semi-honest, meaning that while they may attempt to infer private information, they still adhere to the protocol. (2) At least one of the aggregation servers is trusted, while the other is honest-but-curious. They do not collude with one another or with clients, and together they jointly perform MPC-based aggregation. (3) The model owner is trusted to manage protocol coordination but cannot access client data or intermediate model states. To avoid collusion and information leakage, the trusted server is responsible for pairing clients using the available metadata and the risk-aware dynamic client grouping algorithm. All communication between clients in a group passes through the trusted server to maintain anonymity. The risk-aware dynamic client grouping algorithm further prevents repeated interactions by reassigning clients into new training groups in each round.
Threat Model: We consider a semi-honest (honest-but-curious) adversary who follows the protocol but may attempt to infer private information from observed data. The system involves three types of participants: clients, aggregation servers, and a model owner. Clients perform local training using secret-shared model parameters. Aggregation servers coordinate secure aggregation without learning the underlying values. The model owner is responsible for initializing the global model and orchestrating training (Fig. 2).
Figure 2: Overview of the threat model. Each party retains plaintext within its trusted boundary. Adversaries attempt to infer private information outside the untrusted boundary.
@@
These guarantees mitigate risks such as unauthorized observation, gradient leakage, and reconstruction attacks. Since the model is never exposed in plaintext outside the trusted boundary, no untrusted party can access it to carry out membership inference or model inversion attacks aimed at extracting information about the training user data.
Out-of-Scope Threats: FuSeFL does not currently defend against poisoning attacks [33] and malicious participants who deviate from the protocol, submit corrupted updates, or act in a Byzantine manner. Future extensions may incorporate robust aggregation or verifiable computation to address these threats. Side-channel attacks[34, 35, 36] such as timing[34] or traffic analysis[36] are out of scope and would require separate system-level countermeasures.